OpenStudy (anonymous):

Solve the system of equations by graphing. 2x-y=-2, 4x+y=-4

OpenStudy (lopus):

2x-y=-2 4x+y=-4 ---------- 6x=-6 x=-1 then y=1

OpenStudy (lopus):

excusme y=0

OpenStudy (anonymous):

i need two other points

OpenStudy (lopus):

only have one solution

OpenStudy (anonymous):

Orange is 2x-y=-2 Purple is 4x+y=-4

OpenStudy (anonymous):

do you know the plot points?

OpenStudy (anonymous):

@RoseDryer

OpenStudy (anonymous):

What do you mean by plot points? Like the y-intercept and stuff?

OpenStudy (anonymous):

yes like where did you plug in those lines on the graph? for the first one and second one

OpenStudy (anonymous):

Well I point them in the equation: y=mx+b (where m=slope and b=y-int.) To get to that from the first equation 2x-y=-2 Subtract 2x on both sides to get -y=-2x-2 Divide by -1 on both sides to get y=2x+2 When you graph it 2 in the y-int. and from there you go up 2 over 1 to get to the next point of the line, For he second equation: 4x+y=-4 Subtract 4x on both sides to get y=-4x-4 When you graph it -4 in the y-int. and from there you go down 4 over 1 to get to the next point of the line,